# A stochastic preconditioned Douglas-Rachford splitting method for saddle-point problems

### v2

@@ 4 Sublinear convergence of stochastic PDR for restricted primal-dual gap
$\mathbb{E}[$ $\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]\leq\sup_{z\in\mathbb{B}}\frac{1}{2K}[\|u^{0}-u\|^{2}_{\mathbf{M}_{p}}+\|\mathbf{y}^{0}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}P^{-1}}]$ (4.30)
$-\mathbb{E}\|\mathbf{y}^{K+1}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}+\mathbb{E}\sum_{k=0}^{K}\frac{1}{K}[\frac{1}{2}\|\mathbf{v}^{k+1}\|^{2}_{\mathbf{w}_{y}\mathbf{P}}-\langle\mathbf{y}_{\dagger}^{k},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}].$
@@
Dropping $-\mathbb{E}\|\mathbf{y}^{K+1}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}$, noting $\mathbb{E}\sum_{k=0}^{K}\frac{1}{2}\|\mathbf{v}^{k+1}\|^{2}_{\mathbf{w}_{y}\mathbf{P}}\leq\mathbb{E}[\sum_{k=0}^{K}\frac{1}{2}\|\mathbf{y}^{k}-\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}]$ by (4.15), together with (4.16) and (4.1), we can obtain
@@
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]\leq\frac{1}{2K}[\sup_{z\in\mathbb{B}}(\|u^{0}-u\|^{2}_{\mathbf{M}_{p}}+\|\mathbf{y}^{0}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}})+\Delta^{0}].$ (4.31)
@@
Second, let us give the upper bounds for $\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]$. Denote by $v_{\operatorname{test}}^{k+1}=y_{\operatorname{test}}^{k}-\tilde{y}_{\operatorname{test}}^{k+1}-\mathbf{P}^{-1}(y_{\operatorname{test}}^{k}-y_{\operatorname{test}}^{k+1})$, $y^{k+1}_{\dagger,\operatorname{test}}=y^{k}_{\dagger,\operatorname{test}}-\mathbf{P}v_{\operatorname{test}}^{k+1}$ and select a $y^{0}_{\dagger,\operatorname{test}}\in\mathbb{Y}$ randomly. By the definition of $y^{k}_{\dagger,\operatorname{test}}$, $v^{k}_{\operatorname{test}}$ and (4.14), we obtain
@@
$=$ $\frac{1}{2}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}-\frac{1}{2}\|{y}_{\dagger,\operatorname{test}}^{K+1}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\sum_{k=0}^{K}\frac{1}{2}\|v_{\operatorname{test}}^{k+1}\|_{\mathbf{P}}^{2}-\sum_{k=0}^{K}\langle y^{k}_{\dagger,\operatorname{test}},v^{k+1}_{\operatorname{test}}\rangle$
@@
$\langle\mathcal{K}x,(\mathbf{P}^{-1}-I)(y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0})\rangle\leq\frac{1}{2}\|\mathcal{K}x\|^{2}+\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2},$ (4.33)
@@
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]$ $\leq\sup_{z\in\mathbb{B}}[\frac{1}{2K}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\frac{1}{2K}\|\mathcal{K}x\|^{2}]$
$+\frac{1}{2K}\mathbb{E}\sum_{k=0}^{K}[\|v_{\operatorname{test}}^{k+1}\|_{\mathbf{P}}^{2}+\mathbb{E}\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}K}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2}.$ (4.34)
@@
$\mathbb{E}\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}K}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2}$ $\leq\mathbb{E}\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K}\|y_{\operatorname{test}}^{K+1}-y^{*}\|^{2}+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K}\|y^{*}-y_{\operatorname{test}}^{0}\|^{2}$
$\leq\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K}(n(3C+3\tau)\Delta^{0}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2}).$ (4.35)
@@
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]\leq\sup_{z\in\mathbb{B}}[\frac{1}{2K}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\frac{1}{2K}\|\mathcal{K}x\|^{2}]$
$+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K}(n(3C+3\tau)\Delta^{0}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})+\frac{1}{2\underline{p}K}M_{K}.$ (4.36)
Third, we show $\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))]$ is also bounded. In fact,
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))]=\mathbb{E}\frac{1}{K}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})].$ (4.37)
@@
$\mathbb{E}\frac{1}{K}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})]=\frac{1}{K}\sum_{k=0}^{K}\mathbb{E}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathbb{E}^{k}\mathcal{G}_{\mathbf{P}^{-1}}(y^{k+1}_{\operatorname{test}})+\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k}_{\operatorname{test}})]$
$=\frac{1}{K}\sum_{k=0}^{K}[-\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k+1}_{\operatorname{test}})+\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k}_{\operatorname{test}})]=\frac{1}{K}(\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{0}_{\operatorname{test}})-\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}k}+1}_{\operatorname{test}})).$ (4.38)
@@
$\geq\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)(\mathcal{G}_{i}(y_{i}^{*})-\|\mathcal{K}_{i}x^{*}\|\sqrt{(3C+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test},i}-y^{*}_{i}\|^{2}}),$ (4.39)
@@
$\mathbb{E}\frac{1}{K}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})]$ $\leq\frac{1}{K}\Big{(}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{0}_{\operatorname{test}})-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})$
$+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{(3C+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test},i}-y^{*}_{i}\|^{2}}\Big{)}.$ (4.40)
@@
Finally, with (4.21) and combining (4.31), (4.36) and (4.40), we get (4.19) with the constant $C_{\mathbb{B}}$. For the Bregman distance, replacing $z=(x,y)$ by $(x^{*},y^{*})$, one can obtain the same ergodic convergence rate of the ergodic sequences with similar derivation. ∎
5 Stochastic PDR for quadratic problems (SPDRQ)

### v3

@@ 4 Sublinear convergence of stochastic RPDR for restricted primal-dual gap
$\mathbb{E}[$ $\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]\leq\sup_{z\in\mathbb{B}}\frac{1}{2K_{1}}[\|u^{0}-u\|^{2}_{\mathbf{M}_{p,\varrho_{0}}}+\|\mathbf{y}^{0}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{0}}]$ (4.30)
$-\mathbb{E}\|\mathbf{y}^{K+1}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}+\mathbb{E}\sum_{k=0}^{K}\frac{1}{K_{1}}[\frac{1}{2}\|\mathbf{v}^{k+1}\|^{2}_{\mathbf{w}_{y}\mathbf{P}\mathbf{R}_{y}^{k}}-\langle\mathbf{y}_{\dagger}^{k},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}].$
By (4.15), non-decreasing of $\rho_{k}$, $\rho_{k,y}$ and (4.1) we have,
$\mathbb{E}\sum_{k=0}^{K}\|\mathbf{v}^{k+1}\|^{2}_{\mathbf{w}_{y}\mathbf{P}\mathbf{R}_{y}^{k}}\leq\mathbb{E}[\sum_{k=0}^{K}\|\mathbf{y}^{k}-\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}]\leq\mathbb{E}[\sum_{k=0}^{K}\frac{1}{\rho_{l}}\|\mathbf{y}^{k}-\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}]$ (4.31)
$\leq\mathbb{E}[\sum_{k=0}^{K}\frac{1}{\rho_{l}}\|u^{k}-u^{k+1}\|_{\mathbf{M}_{p}}^{2}]\leq\frac{\rho_{u}^{2}}{(2-\rho_{u})\rho_{l}}\Delta^{0}=\frac{\rho_{u}^{2}}{(2-\rho_{u})\rho_{l}^{2}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}.$
@@
Thus in (4.30), dropping $-\mathbb{E}\|\mathbf{y}^{K+1}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}$, together with (4.16) and (4.31), we can obtain
@@
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]\leq\frac{1}{2K_{1}\rho_{l}}\big{[}\sup_{z\in\mathbb{B}}(\|u^{0}-u\|^{2}_{\mathbf{M}_{p}}+\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}$ (4.32)
$+\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}})+\frac{\rho_{u}^{2}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}\big{]}.$
@@
Second, let us give the upper bounds for $\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]$. Denote by $v_{\operatorname{test}}^{k+1}=y_{\operatorname{test}}^{k}-\tilde{y}_{\operatorname{test}}^{k+1}-\mathbf{P}^{-1}(y_{\operatorname{test}}^{k}-y_{\operatorname{test}}^{k+1})$, $y^{k+1}_{\dagger,\operatorname{test}}=y^{k}_{\dagger,\operatorname{test}}-\mathbf{P}v_{\operatorname{test}}^{k+1}$ and select a $y^{0}_{\dagger,\operatorname{test}}\in\mathbb{B}_{2}$ randomly. By the definition of $y^{k}_{\dagger,\operatorname{test}}$, $v^{k}_{\operatorname{test}}$ and (4.14), we obtain
@@
$\leq$ $\frac{1}{2}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}-\frac{1}{2}\|{y}_{\dagger,\operatorname{test}}^{K+1}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\sum_{k=0}^{K}\frac{1}{2}\|v_{\operatorname{test}}^{k+1}\|_{\mathbf{P}}^{2}-\sum_{k=0}^{K}\langle y^{k}_{\dagger,\operatorname{test}},v^{k+1}_{\operatorname{test}}\rangle$
@@
$\langle\mathcal{K}x,(\mathbf{P}^{-1}-I)(y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0})\rangle\leq\frac{1}{2}\|\mathcal{K}x\|^{2}+\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2}.$ (4.34)
@@
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]$ $\leq\sup_{z\in\mathbb{B}}[\frac{1}{2K_{1}}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\frac{1}{2K_{1}}\|\mathcal{K}x\|^{2}]$
$+\frac{1}{2K_{1}}\mathbb{E}\sum_{k=0}^{K}\|v_{\operatorname{test}}^{k+1}\|_{\mathbf{P}}^{2}+\mathbb{E}\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}K_{1}}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2}.$ (4.35)
@@
$\mathbb{E}\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}K_{1}}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2}$ $\leq\mathbb{E}\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K_{1}}\|y_{\operatorname{test}}^{K+1}-y^{*}\|^{2}+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K_{1}}\|y^{*}-y_{\operatorname{test}}^{0}\|^{2}$
$\leq\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K_{1}}(\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2}).$ (4.36)
@@
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]\leq\sup_{z\in\mathbb{B}}[\frac{1}{2K_{1}}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\frac{1}{2K_{1}}\|\mathcal{K}x\|^{2}]$
$+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K_{1}}(\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})+\frac{1}{2\underline{p}K_{1}}M_{K}.$ (4.37)
Third, we show $\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))]$ is also bounded. In fact,
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))]=\mathbb{E}\frac{1}{K_{1}}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})].$ (4.38)
@@
$\mathbb{E}\frac{1}{K_{1}}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})]=\frac{1}{K_{1}}\sum_{k=0}^{K}\mathbb{E}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathbb{E}^{k}\mathcal{G}_{\mathbf{P}^{-1}}(y^{k+1}_{\operatorname{test}})+\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k}_{\operatorname{test}})]$
$=\frac{1}{K_{1}}\sum_{k=0}^{K}[-\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k+1}_{\operatorname{test}})+\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k}_{\operatorname{test}})]=\frac{1}{K_{1}}(\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{0}_{\operatorname{test}})-\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{{K}+1}_{\operatorname{test}})).$ (4.39)
@@
$\geq\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)(\mathcal{G}_{i}(y_{i}^{*})-\|\mathcal{K}_{i}x^{*}\|\sqrt{\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test},i}-y^{*}_{i}\|^{2}}),$ (4.40)
@@
$\mathbb{E}\frac{1}{K_{1}}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})]$ $\leq\frac{1}{K_{1}}\Big{(}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{0}_{\operatorname{test}})-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})$
$+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test},i}-y^{*}_{i}\|^{2}}\Big{)}.$ (4.41)
@@
Finally, with (4.21) and combining (4.32), (4.37) and (4.41), we get (4.19) with the constant $C_{\mathbb{B}}$. For the Bregman distance, replacing $z=(x,y)$ by $(x^{*},y^{*})$, one can obtain the same sublinear convergence rate for the ergodic sequences. ∎
In the next theorem, we investigate the convergence rate of expected primal error when Lipschitz smoothness is satisfied for $\mathcal{G}$.
Theorem 4.
For the ergodic sequence $z_{\operatorname{test},K}=(x_{\operatorname{test},K},y_{\operatorname{test},K})$ defined in Theorem 3, $K_{1}=K+1$ and any $z=(x,y)\in\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$, if $\mathcal{G}^{*}$ is $\beta$-Lipschitz continuous with respect to the norm $\|\cdot\|_{\frac{I}{\tau}}$, and $y^{0}$, $(\bar{y}^{0}+\tau\mathcal{K}x^{*})$, $y^{0}_{\dagger}$, $(\bar{y}^{0}_{\dagger}+\tau\mathcal{K}x^{*})\in\operatorname{dom}\mathcal{G}$, and if $y^{*}_{\dagger}\in\operatorname{dom}\mathcal{G}$, such that $\sigma\mathcal{K}^{*}y^{*}_{\dagger}=\bar{x}^{0}-x^{*}$, then the expected primal error also exhibits a convergence rate of $\mathcal{O}(1/K)$, i.e,
$\mathbb{E}[\mathcal{F}(x_{\operatorname{test},K})+\mathcal{G}^{*}(\mathcal{K}x_{\operatorname{test},K})-\mathcal{F}(x^{*})-\mathcal{G}^{*}(\mathcal{K}x^{*})]\leq\frac{C_{p,2}}{K_{1}}=\mathcal{O}(1/K),$ (4.42)
where
$C_{p,2}=C_{p}+\frac{1}{2\rho_{l}}(\|x^{0}-x^{*}\|_{\frac{N_{1}-I}{\sigma}}^{2}+4\sigma\tau\beta^{2}\|\mathcal{K}^{*}\|^{2}+8\|N_{2}-I\|\beta^{2}/\underline{p}+8\beta^{2}/\underline{p})$
$\quad\qquad+\frac{1}{2}(\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x^{*}\|_{\mathbf{P}^{-1}}^{2}+\|\mathcal{K}x^{*}\|^{2}),\ \text{with}$
$C_{p}:=\frac{\rho_{u}^{2}}{2(2-\rho_{u})\rho^{2}_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}}(\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})+\frac{1}{2\underline{p}}M_{K}$
$+\mathcal{G}_{\mathbf{P}^{{-1}}-I}(y^{0}_{\operatorname{test}})-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2}}.$
Proof.
Let us first define a new distance function ¯$\bar{\mathcal{H}}(z_{\operatorname{test},K},z)$,
¯$\bar{\mathcal{H}}(z_{\operatorname{test},K},z):=\mathcal{H}(z_{\operatorname{test},K},z)-\frac{1}{2\rho_{l}K_{1}}(\|x^{0}-x\|_{\frac{N_{1}-I}{\sigma}}^{2}+\|y^{0}-y\|_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}^{2}+\|\bar{x}^{0}-x-\sigma\mathcal{K}^{*}y\|^{2}_{\frac{I}{\sigma}}$
$+\|\bar{y}^{0}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}+\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}+\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}})-\frac{1}{2K_{1}}(\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\|\mathcal{K}x\|^{2}),$
where $y^{0}_{\dagger}$, $\bar{y}^{0}_{\dagger}$ and ${y}_{\dagger,\operatorname{test}}^{0}$ are given as in Theorem 3. Following the same proof of Theorem 3, we can drive, $\mathbb{E}\sup_{z\in\mathbb{B}}\bar{\mathcal{H}}(z_{\operatorname{test},K},z)\leq\frac{C_{p}}{K_{1}}$. Then substituting $z=(x^{*},y)$ for any $y\in\mathbb{B}_{2}$ in ¯$\bar{\mathcal{H}}(z_{\operatorname{test},K},z)$, we derive $\mathbb{E}\sup_{y\in\mathbb{B}_{2}}\bar{\mathcal{H}}(z_{\operatorname{test},K},(x^{*},y))\leq\frac{C_{p}}{K_{1}}$. By the definition of $y^{*}_{\dagger}$, we arrive at $\|\bar{x}^{0}-x^{*}-\sigma\mathcal{K}^{*}y\|^{2}_{\frac{I}{\sigma}}\leq\sigma\|\mathcal{K}^{*}\|^{2}\|y^{*}_{\dagger}-y\|^{2}$. With [4, Corollary 17.19] and noting that $\mathcal{G}^{*}_{i}$ is $\beta$-Lipschitz continuous on the norm $\|\cdot\|_{\frac{I}{\tau}}$, we have $\|y^{*}_{\dagger}-y\|^{2}\leq 4\tau\beta^{2}$, $\|y^{0}-y\|_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}^{2}\leq 4\|N_{2}-I\|\beta^{2}/\underline{p}$, $\|\bar{y}^{0}-y+\tau\mathcal{K}x^{*}\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}\leq 4\beta^{2}/\underline{p}$, $\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}\leq 4\|N_{2}-I\|\beta^{2}/\underline{p}$ and $\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x^{*}\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}\leq 4\beta^{2}/\underline{p}$. We thus obtain
$\mathbb{E}\sup_{y\in\mathbb{B}_{2}}\mathcal{H}(z_{\operatorname{test},K},(x^{*},y))$
$=\mathbb{E}\sup_{y\in\mathbb{B}_{2}}[\mathcal{F}(x_{\operatorname{test},K})+\langle\mathcal{K}x_{\operatorname{test},K},y\rangle-\mathcal{G}(y)-\mathcal{F}(x^{*})-\langle\mathcal{K}x^{*},y_{\operatorname{test},K}\rangle+\mathcal{G}(y_{\operatorname{test},K})]\leq\frac{C_{p,2}}{K_{1}}.$ (4.43)
By Lipschitzness of $\mathcal{G}^{*}$, we pick $y\in\partial\mathcal{G}^{*}(\mathcal{K}x_{\operatorname{test},K})\neq\mathbf{0}$ s.t., $\langle\mathcal{K}x_{\operatorname{test},K},y\rangle-\mathcal{G}(y)=\mathcal{G}^{*}(\mathcal{K}x_{\operatorname{test},K})$, and by Fenchel-Young inequality, $\mathcal{G}(y_{\operatorname{test},K})-\langle\mathcal{K}x^{*},y_{\operatorname{test},K}\rangle\geq-\mathcal{G}^{*}(\mathcal{K}x^{*})$. Employing these inequalities in (4.43), we derive (4.42). ∎
5 Stochastic PDR for quadratic problems (SRPDRQ)